# Mining Limited Data Sufficiently: A BERT-inspired Approach for CSI Time Series Application in Wireless Communication and Sensing

### v1

@@ IV-B2 WiFall Dataset
The WiFall dataset is an action recognition, fall detection, and people identification dataset, containing the CSI collected from 10 individuals performing 5 different actions, including walking, jumping, sitting, standing up, and falling, which also consists of many types like forward fall, left fall, right fall, seated fall, and backward fall.
@@ IV-B3 WiCount Dataset
The WiCount dataset is a people number estimation dataset, collected in a small meeting room as shown in Fig. 5. The sketch map of the dataset is illustrated as Fig. 6 that includes people number from 0 to 3. When collecting the data, people in the environment moved casually that they may jump, squat down, walk forward or backward, sit, and so on.
Figure 5: Data Collection Environment of WiCount Dataset: The unit in the picture is meters.
@@
Figure 6: Sketch Map of WiCount Dataset
@@ IV-C Experiment Result
IV-C1 Pre-training Result
In this section, we first illustrate the performance of our model during the pre-training phase, as this directly influences the model’s performance and convergence speed in downstream tasks. Following the approach taken by most signal generation studies [44], we evaluate pre-training performance using two metrics.
The first dimension is the signal recovery error. To demonstrate that CSI-BERT2 effectively learns the CSI pattern, we compare the recovery performance of our method against that of CSI-BERT1 and various traditional interpolation methods. Specifically, we assess the recovery performance by randomly deleting 15% of the packets and comparing the model’s recovery results with the ground truth. We employ Mean Squared Error (MSE), Symmetric Mean Absolute Percentage Error (SMAPE), and Mean Absolute Percentage Error (MAPE) to quantify the error, as represented in Eq. 12:
$\text{MSE}(\hat{c_{i}},c_{i})=\frac{1}{n}\sum_{i=1}^{n}(c_{i}-\hat{c_{i}})^{2}\ ,$ (12)
$\text{SMAPE}(\hat{c_{i}},c_{i})=\frac{1}{n}\sum_{i=1}^{n}\frac{|c_{i}-\hat{c_{i}}|}{(\frac{|c_{i}|+|\hat{c_{i}}|}{2}+\epsilon)}\ ,$
$\text{MAPE}(\hat{c_{i}},c_{i})=\frac{1}{n}\sum_{i=1}^{n}\frac{|c_{i}-\hat{c_{i}}|}{c_{i}+\epsilon}\ ,$
where $c_{i}$ and $\hat{c_{i}}$ represent ground truth and model output respectively, $n$ represents the number of samples, $\epsilon$ is a small number used to avoid division by zero. The metrics are calculated only on the 15% of the deleted CSI. Furthermore, to ensure fairness, we compute the error solely on the testing set, rather than on the entire dataset as done in CSI-BERT1. We also compare the time required to recover the full dataset across different models, using only the CPU for this comparison. Our results indicate that CSI-BERT2 achieves relatively low recovery error and recovery time across most datasets.
@@ IV-C1 Pre-training Result
Dataset WiGesture [24] WiFall[26] WiCount Method Metric MSE SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s) CSI-BERT2 2.0800 0.1153 0.1217 5.53 4.1463 0.1240 0.1351 2.77 2.4531 0.1092 0.1189 1.56 CSI-BERT1 [24] 2.2438 0.1156 0.1244 1.84 4.4042 0.1271 0.1373 1.32 2.4471 0.1092 0.1185 0.67 Linear Interpolation 2.8642 0.1266 0.1364 38.49 6.4420 0.1461 0.1571 2.81 2.6870 0.1099 0.1175 1.20 Ordinary Kringing 3.5090 0.1390 0.1612 2709.43 4.6637 0.1319 0.1462 289.09 4.5964 0.1423 0.1684 109.10 Inverse Distance Weighted (IDW) 2.4726 0.1187 0.1301 19.82 4.4251 0.1276 0.1409 2.45 3.4431 0.1268 0.1483 0.82
However, a low recovery error does not necessarily imply that the model has effectively learned the CSI pattern. For instance, the vanilla BERT model [20] also achieves a relatively low recovery error by producing similar outputs regardless of the input. Consequently, we adopt a common approach in generation studies by using the recovered data to train other models. If these models perform well on the generated data, it suggests that the recovered data is likely to be close to reality. Following the methodology of CSI-BERT1 [24], we provide two methods for CSI recovery, named “recover” and “replace”:
@@
Subsequently, we utilized the data recovered by the different methods to train various networks and evaluate their accuracy across multiple tasks, including gesture recognition, person identification, action recognition, fall detection, and crowd estimation. The models compared include traditional neural networks such as MLP, recently proposed architectures like LSTM [45], and Wi-Fi sensing models such as WiGRUNT [46]. The results are summarized in Table III. The findings clearly indicate that using CSI-BERT2 for data recovery yields the most significant improvements compared to other methods, suggesting that the data recovered by CSI-BERT2 is the closest to reality. Through these two aspects, we demonstrate the effectiveness of our pre-training method.
@@
Task Gesture Recognition (WiGesture Dataset [24]) Data Model MLP CNN RNN [47] LSTM [45] Chen et al. [48] WiGRUNT [46] CSI-BERT1 [24] CSI-BERT2 Average 337K 23K 33K 133K 11M 11M 2M 5M Original Data 66.93% 55.72% 39.56% 11.97% 70.31% 48.73% 76.91% 99.48% 58.70% CSI-BERT2 recover 72.88% 57.27% 54.34% 48.35% 92.96% 78.97% 92.18% 89.06% 73.25% CSI-BERT2 replace 73.68% 62.80% 55.48% 40.79% 91.92% 74.99% 81.51% 91.95% 71.63% CSI-BERT1 recover 74.23% 59.39% 48.96% 22.92% 92.57% 71.87% 71.87% 92.70% 66.81% CSI-BERT1 replace 86.90% 61.51% 58.80% 52.36% 84.52% 78.84% 79.54% 91.41% 74.24% Linear Interpolation 72.91% 58.35% 45.32% 49.09% 80.75% 74.91% 74.55% 88.25% 68.01% Ordinary Kringing 65.62% 57.55% 53.64% 50.00% 88.71% 69.99% 74.27% 85.93% 68.21% IDW 40.17% 56.77% 48.70% 46.88% 80.32% 71.06% 67.22% 88.28% 62.42% Task People Identification (WiGesture Dataset [24]) Model MLP CNN RNN [47] LSTM [45] Chen et al. [48] WiGRUNT [46] CSI-BERT1 [24] CSI-BERT2 Average Original Data 71.34% 71.14% 66.39% 21.09% 83.76% 72.07% 93.94% 99.73% 72.43% CSI-BERT2 recover 95.57% 85.54% 84.60% 27.98% 93.20% 81.73% 97.92% 99.73% 83.28% CSI-BERT2 replace 95.05% 83.07% 84.68% 54.13% 95.33% 83.84% 96.35% 94.79% 85.91% CSI-BERT1 recover 97.13% 80.60% 80.51% 35.18% 94.30% 84.67% 95.05% 99.73% 83.39% CSI-BERT1 replace 97.65% 79.18% 89.24% 24.22% 97.39% 77.77% 95.83% 99.47% 82.59% Linear Interpolation 81.84% 70.88% 84.45% 26.83% 86.75% 70.28% 97.92% 91.67% 76.33% Ordinary Kringing 94.76% 85.38% 86.42% 21.61% 97.32% 80.84% 95.83% 99.03% 82.64% IDW 83.22% 74.56% 88.54% 33.91% 94.27% 80.70% 95.20% 99.47% 81.23% Task Action Recognition (WiFall Dataset [26]) Model MLP CNN RNN [47] LSTM [45] Chen et al. [48] WiGRUNT [46] CSI-BERT1 [24] CSI-BERT2 Average Original Data 47.48% 56.27% 58.61% 52.10% 51.38% 34.44% 82.43% 88.59% 58.91% CSI-BERT2 recover 64.97% 67.18% 68.48% 62.63% 71.70% 70.96% 67.63% 72.16% 68.21% CSI-BERT2 replace 69.01% 66.27% 70.18% 61.99% 73.96% 69.72% 66.77% 72.70% 68.82% CSI-BERT1 recover 66.40% 54.94% 68.48% 61.79% 69.66% 70.94% 67.36% 73.69% 66.65% CSI-BERT1 replace 73.05% 54.97% 66.79% 66.73% 72.01% 67.44% 66.61% 73.67% 67.65% Linear Interpolation 67.44% 64.32% 67.31% 59.78% 74.22% 70.57% 64.37% 74.19% 67.77% Ordinary Kringing 67.96% 65.52% 64.44% 63.88% 70.92% 62.41% 67.36% 71.77% 66.78% IDW 70.31% 67.08% 69.79% 62.32% 71.09% 72.39% 67.22% 70.21% 68.80% Task Fall Detection (WiFall Dataset [26]) Model MLP CNN RNN [47] LSTM [45] Chen et al. [48] WiGRUNT [46] CSI-BERT1 [24] CSI-BERT2 Average Original Data 78.34% 52.99% 82.29% 80.35% 78.52% 73.69% 93.28% 94.79% 79.28% CSI-BERT2 recover 80.79% 75.95% 86.58% 86.72% 82.42% 80.90% 82.25% 86.97% 82.82% CSI-BERT2 replace 79.82% 74.89% 83.07% 86.31% 82.16% 78.65% 80.62% 85.38% 81.36% CSI-BERT1 recover 80.98% 75.27% 84.37% 80.41% 81.38% 83.07% 81.32% 84.92% 81.46% CSI-BERT1 replace 80.21% 74.94% 83.46% 84.37% 82.33% 80.79% 83.33% 85.72% 81.89% Linear Interpolation 81.78% 75.78% 84.50% 84.33% 78.51% 78.77% 81.35% 84.39% 81.17% Ordinary Kringing 81.64% 75.78% 80.98% 82.29% 82.00% 79.03% 82.31% 84.49% 81.07% IDW 82.55% 54.94% 83.59% 80.59% 78.21% 80.72% 81.72% 84.06% 78.29% Task People Number Estimation (WiCount Dataset) Model MLP CNN RNN [47] LSTM [45] Chen et al. [48] WiGRUNT [46] CSI-BERT1 [24] CSI-BERT2 Average Original Data 56.77% 69.68% 80.93% 80.72% 48.33% 49.53% 89.67% 94.32% 71.24% CSI-BERT2 recover 87.29% 78.75% 83.98% 81.51% 83.32% 85.42% 84.06% 91.32% 84.45% CSI-BERT2 replace 85.62% 78.49% 88.12% 86.97% 81.41% 82.34% 81.51% 92.76% 84.65% CSI-BERT1 recover 88.98% 80.83% 86.20% 82.39% 82.22% 82.70% 79.04% 92.70% 84.38% CSI-BERT1 replace 81.61% 72.60% 85.67% 84.95% 85.62% 82.75% 81.61% 92.86% 83.46% Linear Interpolation 76.51% 77.73% 85.52% 82.40% 80.17% 83.07% 86.51% 88.64% 82.57% Ordinary Kringing 87.29% 50.52% 84.84% 85.05% 82.97% 76.72% 85.72% 91.90% 80.63% IDW 80.72% 82.29% 84.17% 87.00% 82.10% 81.72% 85.62% 88.54% 84.02%
@@ IV-C2 CSI Prediction Task
Due to the scarcity of real CSI data in wireless communication, we follow the approach taken by most previous studies and utilize wireless sensing datasets to evaluate model performance. For the prediction task, we provide the first 0.8 seconds of CSI in each sample and train the model to predict the subsequent 0.2 seconds of CSI. We compare the performance of our CSI-BERT2 model against several other CSI prediction models, including popular time series models like LSTM [45], RNN [47], GRU [49], and Mamba [50], which are also widely used in CSI prediction [51, 52, 53, 54, 55, 56]. We also compare our model with some CSI prediction networks like OCEAN [57], CV-3DCNN [58], and ConvLSTM [59]. The prediction error and time cost (using GPU) are shown in Table IV. It can be seen that our CSI-BERT2 model significantly outperforms all other models. Particularly in the WiCount dataset, where the task is more challenging, as each gesture or action has a relatively irregular pattern, other methods exhibit a substantial decrease in performance, while CSI-BERT2 is influenced only minimally. Although the inference time of CSI-BERT2 is longer than that of the other models, its prediction time for the entire testing set across different datasets remains under 0.5 seconds, which is acceptable.
@@
Dataset WiGesture [24] WiFall[26] WiCount Method Metric MSE SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s) CSI-BERT2 (5M) 3.2942 0.1583 0.1349 0.46 4.8598 0.1471 0.1347 0.49 5.3401 0.1726 0.1590 0.46 LSTM (133K) [45, 51, 52, 53, 54] 12.3254 0.2397 0.3967 0.05 7.1495 0.1624 0.1882 0.04 32.3377 0.2547 0.3528 0.05 RNN (33K) [47, 55] 19.4708 0.2877 0.4063 0.04 16.9083 0.2424 0.2988 0.04 32.3670 0.2548 0.3534 0.03 GRU (100K) [49, 56] 19.7180 0.2922 0.4243 0.04 16.5353 0.2395 0.2963 0.07 39.8108 0.2541 0.3556 0.02 Mamba (5M) [50] 12.3281 0.2392 0.3277 0.24 6.4666 0.1532 0.1756 0.12 39.9170 0.2566 0.3524 0.11 OCEAN (126K) [57] 19.6257 0.2925 0.4231 0.05 16.8825 0.2423 0.2978 0.03 39.7917 0.2542 0.3548 0.02 CV-3DCNN (19K) [58] 11.3017 0.2267 0.3044 0.04 8.2616 0.1713 0.1981 0.03 42.2662 0.2631 0.3560 0.02 ConvLSTM (152K) [59] 19.7038 0.2921 0.4242 0.04 16.8935 0.2429 0.2983 0.03 39.7709 0.2537 0.3552 0.02

### v2

@@ IV-B2 WiFall Dataset
The WiFall dataset is an action recognition, fall detection, and people identification dataset, containing the CSI collected from 10 individuals performing 5 different actions, including walking, jumping, sitting, standing up, and falling, which also consists of various types like forward fall, left fall, right fall, seated fall, and backward fall.
@@ IV-B3 WiCount Dataset
The WiCount dataset is a people number estimation dataset, collected in a small meeting room as shown in Fig. 4. The sketch map of the dataset is illustrated in Fig. 5 that includes people number from 0 to 3. During data collection, individuals moved naturally, performing various activities such as walking, squatting, jumping, sitting, and moving in different directions.
Figure 4: Data collection environment of WiCount dataset, with measurements shown in meters.
@@
Figure 5: Sketch map of WiCount dataset.
IV-B4 CommPre Dataset
The CommPre dataset is generated using the Sionna RT[46] platform with ray tracing techniques that capture multiple electromagnetic phenomena, including reflection, refraction, and diffraction. To realistically simulate challenging high-mobility environments, the positions of objects within the environment dynamically vary over time, with velocities ranging from 5 m/s to 20 m/s. Diverse mobility patterns are configured across different scenarios to represent various movement dynamics, and the corresponding multipath gains and phases are computed accordingly. Subsequently, the CSI is derived following Eq. (1), producing a CSI matrix of size 100 × 52 for each of the 892 simulated scenarios.
@@ IV-C Experiment Result
IV-C1 Unsupervised Training Result
In this section, we first illustrate the performance of our model during the unsupervised training phase, as this directly influences the model’s performance and convergence speed in downstream tasks. Following the approach taken by most signal generation studies [47], we evaluate unsupervised training performance using two metrics, including signal recovery error and the performance of the downstream model trained on the recovered signal.
The first dimension is the signal recovery error. To demonstrate that CSI-BERT2 effectively learns the CSI pattern, we compare the recovery performance of our method against that of CSI-BERT and other traditional interpolation methods. Specifically, we assess the recovery performance by randomly deleting 15% of the packets and comparing the model’s recovery results with the ground truth. We employ mean squared error (MSE), symmetric mean absolute percentage error (SMAPE), and mean absolute percentage error (MAPE) to quantify the error:
$\text{MSE}(\hat{c},c)=\frac{1}{NM}\sum_{n=1}^{N}\sum_{m=1}^{M}(c^{(m)}_{n}-\hat{c}^{(m)}_{n})^{2}\ ,$ (11)
$\text{SMAPE}(\hat{c},c)=\frac{1}{NM}\sum_{n=1}^{N}\frac{|c^{(m)}_{n}-\hat{c}_{n}^{(m)}|}{(\frac{|c^{(m)}_{n}|+|\hat{c}^{(m)}_{n}|}{2}+\epsilon)}\ ,$
$\text{MAPE}(\hat{c},c)=\frac{1}{NM}\sum_{n=1}^{N}\frac{|c^{(m)}_{n}-\hat{c}^{(m)}_{n}|}{c^{(m)}_{n}+\epsilon}\ ,$
[…]
@@ IV-C1 Unsupervised Training Result
Dataset WiGesture [7] WiFall[37] WiCount Method Metric MSE SMAPE MAPE Time (s) MSE SMAPE MAPE Time(s) MSE SMAPE MAPE Time (s) CSI-BERT2 2.0800 0.1153 0.1217 5.53 4.1463 0.1240 0.1351 2.77 2.4531 0.1092 0.1189 1.56 CSI-BERT [7] 2.2438 0.1156 0.1244 1.84 4.4042 0.1271 0.1373 1.32 2.4471 0.1092 0.1185 0.67 Linear Interpolation 2.8642 0.1266 0.1364 38.49 6.4420 0.1461 0.1571 2.81 2.6870 0.1099 0.1175 1.20 Ordinary Kringing 3.5090 0.1390 0.1612 2709.43 4.6637 0.1319 0.1462 289.09 4.5964 0.1423 0.1684 109.10 Inverse Distance Weighted (IDW) 2.4726 0.1187 0.1301 19.82 4.4251 0.1276 0.1409 2.45 3.4431 0.1268 0.1483 0.82
However, a low recovery error does not necessarily imply that the model has effectively learned the CSI pattern. For instance, the vanilla BERT model [18] also achieves a relatively low recovery errors by producing similar outputs regardless of the input. Consequently, we adopt a common approach in generation studies by using the recovered data to train other models. If these models perform well on the generated data, it suggests that the recovered data is likely to be close to reality. Following the methodology of CSI-BERT [7], we provide two methods for CSI recovery, named “recover” and “replace”:
@@
Subsequently, we utilize the data recovered by different methods to train various networks and evaluate their accuracy across multiple tasks, including gesture recognition, person identification, action recognition, fall detection, and crowd estimation. The models for comparision include traditional neural networks such as MLP, recently proposed architectures like LSTM [48], and Wi-Fi sensing models such as WiGRUNT [49]. The results are summarized in Table III. The findings clearly indicate that using CSI-BERT2 for data recovery yields the most significant improvements compared to other methods, suggesting that the data recovered by CSI-BERT2 is the closest to reality.
@@
Task Gesture Recognition (WiGesture Dataset [7]) Data Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average 337K 23K 33K 133K 11M 11M 2M 5M Original Data 66.93% 55.72% 39.56% 11.97% 70.31% 48.73% 76.91% 99.48% 58.70% CSI-BERT2 recover 72.88% 57.27% 54.34% 48.35% 92.96% 78.97% 92.18% 89.06% 73.25% CSI-BERT2 replace 73.68% 62.80% 55.48% 40.79% 91.92% 74.99% 81.51% 91.95% 71.63% CSI-BERT recover 74.23% 59.39% 48.96% 22.92% 92.57% 71.87% 71.87% 92.70% 66.81% CSI-BERT replace 86.90% 61.51% 58.80% 52.36% 84.52% 78.84% 79.54% 91.41% 74.24% Linear Interpolation 72.91% 58.35% 45.32% 49.09% 80.75% 74.91% 74.55% 88.25% 68.01% Ordinary Kringing 65.62% 57.55% 53.64% 50.00% 88.71% 69.99% 74.27% 85.93% 68.21% IDW 40.17% 56.77% 48.70% 46.88% 80.32% 71.06% 67.22% 88.28% 62.42% Task People Identification (WiGesture Dataset [7]) Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average Original Data 71.34% 71.14% 66.39% 21.09% 83.76% 72.07% 93.94% 99.73% 72.43% CSI-BERT2 recover 95.57% 85.54% 84.60% 27.98% 93.20% 81.73% 97.92% 99.73% 83.28% CSI-BERT2 replace 95.05% 83.07% 84.68% 54.13% 95.33% 83.84% 96.35% 94.79% 85.91% CSI-BERT recover 97.13% 80.60% 80.51% 35.18% 94.30% 84.67% 95.05% 99.73% 83.39% CSI-BERT replace 97.65% 79.18% 89.24% 24.22% 97.39% 77.77% 95.83% 99.47% 82.59% Linear Interpolation 81.84% 70.88% 84.45% 26.83% 86.75% 70.28% 97.92% 91.67% 76.33% Ordinary Kringing 94.76% 85.38% 86.42% 21.61% 97.32% 80.84% 95.83% 99.03% 82.64% IDW 83.22% 74.56% 88.54% 33.91% 94.27% 80.70% 95.20% 99.47% 81.23% Task Action Recognition (WiFall Dataset [37]) Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average Original Data 47.48% 56.27% 58.61% 52.10% 51.38% 34.44% 82.43% 88.59% 58.91% CSI-BERT2 recover 64.97% 67.18% 68.48% 62.63% 71.70% 70.96% 67.63% 72.16% 68.21% CSI-BERT2 replace 69.01% 66.27% 70.18% 61.99% 73.96% 69.72% 66.77% 72.70% 68.82% CSI-BERT recover 66.40% 54.94% 68.48% 61.79% 69.66% 70.94% 67.36% 73.69% 66.65% CSI-BERT replace 73.05% 54.97% 66.79% 66.73% 72.01% 67.44% 66.61% 73.67% 67.65% Linear Interpolation 67.44% 64.32% 67.31% 59.78% 74.22% 70.57% 64.37% 74.19% 67.77% Ordinary Kringing 67.96% 65.52% 64.44% 63.88% 70.92% 62.41% 67.36% 71.77% 66.78% IDW 70.31% 67.08% 69.79% 62.32% 71.09% 72.39% 67.22% 70.21% 68.80% Task Fall Detection (WiFall Dataset [37]) Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average Original Data 78.34% 52.99% 82.29% 80.35% 78.52% 73.69% 93.28% 94.79% 79.28% CSI-BERT2 recover 80.79% 75.95% 86.58% 86.72% 82.42% 80.90% 82.25% 86.97% 82.82% CSI-BERT2 replace 79.82% 74.89% 83.07% 86.31% 82.16% 78.65% 80.62% 85.38% 81.36% CSI-BERT recover 80.98% 75.27% 84.37% 80.41% 81.38% 83.07% 81.32% 84.92% 81.46% CSI-BERT replace 80.21% 74.94% 83.46% 84.37% 82.33% 80.79% 83.33% 85.72% 81.89% Linear Interpolation 81.78% 75.78% 84.50% 84.33% 78.51% 78.77% 81.35% 84.39% 81.17% Ordinary Kringing 81.64% 75.78% 80.98% 82.29% 82.00% 79.03% 82.31% 84.49% 81.07% IDW 82.55% 54.94% 83.59% 80.59% 78.21% 80.72% 81.72% 84.06% 78.29% Task People Number Estimation (WiCount Dataset) Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average Original Data 56.77% 69.68% 80.93% 80.72% 48.33% 49.53% 89.67% 94.32% 71.24% CSI-BERT2 recover 87.29% 78.75% 83.98% 81.51% 83.32% 85.42% 84.06% 91.32% 84.45% CSI-BERT2 replace 85.62% 78.49% 88.12% 86.97% 81.41% 82.34% 81.51% 92.76% 84.65% CSI-BERT recover 88.98% 80.83% 86.20% 82.39% 82.22% 82.70% 79.04% 92.70% 84.38% CSI-BERT replace 81.61% 72.60% 85.67% 84.95% 85.62% 82.75% 81.61% 92.86% 83.46% Linear Interpolation 76.51% 77.73% 85.52% 82.40% 80.17% 83.07% 86.51% 88.64% 82.57% Ordinary Kringing 87.29% 50.52% 84.84% 85.05% 82.97% 76.72% 85.72% 91.90% 80.63% IDW 80.72% 82.29% 84.17% 87.00% 82.10% 81.72% 85.62% 88.54% 84.02%
@@ IV-C2 CSI Prediction Task
Given the scarcity of real CSI datasets in wireless communications, we adopt the widely used approach of leveraging wireless sensing datasets for performance evaluation. For the prediction task, we provide the first 0.8 seconds of CSI in each sample and train the model to predict the subsequent 0.2 seconds of CSI. We compare the performance of our CSI-BERT2 model against several other CSI prediction models, including popular time series models like LSTM [48], RNN [50], GRU [52], and Mamba [53], which are also widely used in CSI prediction [8, 9, 10, 11, NN_CSI, 54]. We also compare our model with some CSI prediction networks like OCEAN [55], CV-3DCNN [56], and ConvLSTM [57]. The prediction error and time cost (using GPU) are shown in Table IV. It can be seen that our CSI-BERT2 model significantly outperforms all other models. Particularly in the WiCount dataset, where the task is more challenging, as each gesture or action has a relatively irregular pattern, other methods exhibit a substantial decrease in performance, while CSI-BERT2 is influenced only minimally. Although the inference time of CSI-BERT2 is longer than that of the other models, its prediction time for the entire testing set across different datasets remains under 0.5 seconds, which is acceptable.
@@
Dataset CommPre WiGesture [7]
Method Metric MSE (×$\times 10^{-8}$) SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s)
CSI-BERT2 (5M) 1.5689 0.2959 0.2319 0.37 3.2942 0.1583 0.1349 0.46
LSTM (133K) [48, 8, 9, 10, 11] 5.1107 0.4849 0.5722 0.03 12.3254 0.2397 0.3967 0.05
RNN (33K) [50, 58] 4.6164 0.4607 0.5385 0.03 19.4708 0.2877 0.4063 0.04
GRU (100K) [52, 54] 5.4009 0.5128 0.6367 0.04 19.7180 0.2922 0.4243 0.04
Mamba (5M) [53] 5.1286 0.4430 0.4838 0.23 12.3281 0.2392 0.3277 0.24
OCEAN (126K) [55] 5.6013 0.4863 0.5576 0.05 19.6257 0.2925 0.4231 0.05
CV-3DCNN (19K) [56] 4.5276 0.4794 0.6810 0.04 11.3017 0.2267 0.3044 0.04
ConvLSTM (152K) [57] 5.0015 0.4898 0.6031 0.04 19.7038 0.2921 0.4242 0.04
Dataset WiFall[37] WiCount
Method Metric MSE SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s)
CSI-BERT2 (5M) 4.8598 0.1471 0.1347 0.49 5.3401 0.1726 0.1590 0.46
LSTM (133K) [48, 8, 9, 10, 11] 7.1495 0.1624 0.1882 0.04 32.3377 0.2547 0.3528 0.05
RNN (33K) [50, 58] 16.9083 0.2424 0.2988 0.04 32.3670 0.2548 0.3534 0.03
GRU (100K) [52, 54] 16.5353 0.2395 0.2963 0.07 39.8108 0.2541 0.3556 0.02
Mamba (5M) [53] 6.4666 0.1532 0.1756 0.12 39.9170 0.2566 0.3524 0.11
OCEAN (126K) [55] 16.8825 0.2423 0.2978 0.03 39.7917 0.2542 0.3548 0.02
CV-3DCNN (19K) [56] 8.2616 0.1713 0.1981 0.03 42.2662 0.2631 0.3560 0.02
ConvLSTM (152K) [57] 16.8935 0.2429 0.2983 0.03 39.7709 0.2537 0.3552 0.02